OpenStudy (anonymous):

A carpenter has a board 4 feet 3 inches in length. He cuts off a piece 2 feet 8 inches in length . The length of the piece that is left is ? Help please

OpenStudy (anonymous):

1 foot and 5 inches

OpenStudy (anonymous):

1 foot = 12 inches so the original length = 4*12 +3 = 51 inches portion cut-off = 2 feet and 8 inches = 2*12 +8 = 32 inches remaining length = ______? divide the remaining length by 12, write the quotient is in feets and the remainder in inches.

OpenStudy (anonymous):

remaining length in inches = 19 inches so in fraction form it would be 1 7/12

OpenStudy (anonymous):

I got 1ft .7 in

OpenStudy (anonymous):

Thank you much